Assume the following for the economy of a​ country: bullet  Consumption​ function: Upper C equals 60 plus 0.75 Upper Y Subscript d bullet  ​Investment: I​ = 90 bullet  Government​ spending: G​ = 56 bullet  Net​ taxes: Upper T equals 0.2 Upper Y minus 25 bullet  Disposable​ income: Upper Y Subscript d Baseline equals Upper Y minus Upper T bullet  ​Equilibrium: Y​ = C​ + I​ + G

What is equilibrium income?
Answer #1

C=60+0.75*Yd....Consumption Function

I=90...Investment Function

G=56....Government Expenditure

T=0.2*Y-25...Taxes and transfers

Yd=Y-T=Y-0.25*Y+25...Disposable Income

Y=C+I+G

=60+0.75((0..75*Y+25)+90+56

Y=60+0.5625*Y-18.75+146

0.4375*Y=187.25

Y=428...Equilibrium INcome